Reserve Problems Chapter 11 Section 2 Problem 2 x Your answer is incorrect. Try again. The peanut crop was harvested from fivUse excel
Answer #1

Install analysis toolpak in excel

Go to Data>Data analysis>Regression

Select Y as Y

X as X

Click ok

Y SUMMARY OUTPUT 2.55 4.18 3.64 4.23 7170 15570 13680 20070 13120 Regression Statistics Multiple R 0.88156 R Square 0.777148

dt SS Regression Residual Total 1 3 MS F Significance F 67460430.18 67460430.18 10.46187 0.048052 19344649.82 6448216.606 868

from output:

ANSWER(a)

common variance=sigma^2=MSE=6448217

ANSWER(B)

beta1=5399

ANSWER(C)

we have

y^=-4942.9+5399.227*4.23

y^= 17895.83

y^= 17896

erro=observed y-predicted y

=20070-17895.83

=2174.17

e=2174

ANSWER(D)

y^=-4942.9+5399.227*2

=5855.554

y^=5856